# Price of Uncertainty for Consensus Games

### v1

@@ 3.2 Proof of Improved Upper Bound
The quantity $\sum E_{0}$ grows on the order of $O\left(m|S_{0}|\right)$.
Proof.
Since there are only $n$ vertices in the graph, every vertex in $S_{0}$ can have at most $n$ edges. Recall that $\sum E_{0}$ counts the number of edges incident to a vertex in $S_{0}$ and a vertex in $G\setminus S_{0}$. By Lemma 8, we have
$\sum E_{0}$ $=O\left(n|S_{0}|\right)$
$\sum E_{0}$ $=O\left(m|S_{0}|\right).$
We will now consider the time reversed process of playing moves on the $E_{k}$.
Definition 12.
A reversed move on a sequence $E_{k}$ consists of the following steps:
  • Select a list of positive elements $\{e_{1},e_{2},\dots,e_{\alpha}\}$ from $E_{k}$
  • Subtract one from each of the $e_{j}$ for $1\leq j\leq\alpha$
  • Calculate $z=\lceil\alpha/(1+\epsilon)\rceil$
  • Create $E_{k-1}$ by inserting $z$ into $E_{k}$ as a new element
Let $\sum E_{k}^{2}$ denote the sum of the squares of the elements of $E_{k}$. In order to investigate the asymptotic growth rate of $\sum E_{m}$, we wish to find the largest starting value of $\sum E_{m}$ for which it is possible to play $m$ reversed moves to get to a desired value of $\sum E_{0}$.
Our plan is to bound the behavior of $\sum E_{k}$ and $\sum E_{k}^{2}$ as reversed moves are played. Intuitively, playing a reversed move on many elements of $E_{k}$ would result in a large decrease in from $\sum E_{k}$ to $\sum E_{k-1}$. However, playing such a move would also result in the creation of a new, very large element in $E_{k-1}$. Since we can only ever subtract one from each element per move, having the sum $\sum E_{k-1}$ concentrated in a few large elements will make decreasing $\sum E$ more difficult. We consider $\sum E_{k}^{2}$ as a way to measure how ”spread out” the sequence $E_{k}$ is, and this particular quantity is preferable over other measures because of how the change from $\sum E_{k}^{2}$ to $\sum E_{k-1}^{2}$ can be cleanly bounded by an expression in terms of $\sum E_{k}$.
Consider some arbitrary sequence $E_{k}$. Let $\alpha$ denote the number of elements of $E$ that we play the current reversed game move on, and let these elements be denoted $e_{1},e_{2},\dots,e_{\alpha}$. By the rules of the reversed game, playing such a move will create a new element of size $z=\lceil\alpha/(1+\varepsilon)\rceil$. It is easy to see that the effect of such a move on $\sum E_{k}$ is
$\sum E_{k-1}=\sum E_{k}-\alpha+z.$
Next, we must have
$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-\sum_{j=1}^{\alpha}\left(e_{j}^{2}-(e_{j}-1)^{2}\right)+z^{2}$
$=\sum E_{k}^{2}-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}.$
Since $e_{1},e_{2},\dots,e_{\alpha}$ are all elements of $E_{k}$, their sum must be bounded by $\sum E_{k}$. By plugging this into our expression for the change in $\sum E^{2}$, we get
$\sum E_{k-1}^{2}-\sum E_{k}^{2}$
$={}$ $-2\sum_{j=1}^{\alpha}e_{j}+\alpha+z^{2}$
$\geq{}$ $-2\sum_{j=1}^{\alpha}e_{j}+z^{2}$
$\geq{}$ $-2\sum E_{k}+\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil^{2}$
$\geq{}$ $-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}$ (1)
Lemma 13.
It is always optimal to take the lower bound in (1) for the change in $\sum E^{2}$.
Proof.
Notice that neither our change in $\sum E$ nor our bound (1) on the change in $\sum E^{2}$ depend on $\sum E_{k}^{2}$. If we make $\sum E_{k}^{2}$ as small as possible for a fixed $\sum E_{k}$, the future options we will have for $\alpha$ will not change. Furthermore, we only care that $\sum E_{0}^{2}$ ends up being smaller than $\left(\sum E_{0}\right)^{2}$. This means that allowing $\sum E_{k-1}^{2}$ to attain its lower bound will never result in more moves being required. There are no further issues because there is no need that intermediate values can be represented by actual sequences; there is no required minimum value for $\sum E_{k}^{2}$ for any integer $0\leq k\leq m-1$. ∎
We can also obtain a simpler lower bound for the change in $\sum E$.
$\sum E_{k-1}-\sum E_{k}$
$={}$ $-\alpha+z$
$={}$ $\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
$\geq{}$ $(1-\varepsilon)\alpha-\alpha$
$={}$ $-\varepsilon\alpha.$ (2)
Lemma 14.
It is always optimal to take the lower bound in (2) for the change in $\sum E$.
Proof.
If we allow $\alpha\in\mathbb{R}$, still with $\alpha\geq 1$, it is always possible to choose a new $\alpha^{\prime}$ such that
$\varepsilon\alpha^{\prime}=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha.$
The RHS of the previous equation is clearly nonpositive, so $\alpha^{\prime}$ must be nonpositive as well. We can also bound the size of $\alpha^{\prime}$ with
$-\varepsilon\alpha^{\prime}$ $=\left\lceil\frac{\alpha}{1+\varepsilon}\right\rceil-\alpha$
$1-\varepsilon\alpha^{\prime}$ $\geq 1+\frac{\alpha}{1+\varepsilon}-\alpha$
$-\varepsilon\alpha^{\prime}$ $\geq\frac{-\varepsilon\alpha}{1+\varepsilon}$
$(1+\varepsilon)\alpha^{\prime}$ $\leq\alpha$
$\alpha^{\prime}$ $<\alpha.$
By (1), if we take the lower bound for the change in $\sum E_{k}^{2}$ using $\alpha^{\prime}$, the value of $\sum E_{k-1}^{2}$ is equal to or smaller than if we had used $\alpha$ instead. By the same logic that was used to prove Lemma 13, it is always helpful to have a lower $\sum E^{2}$. ∎
With Lemma 13 and Lemma 14, we now have a strictly stronger set of actions on $\sum E$ and $\sum E^{2}$ that we can perform with each move.
Definition 15.
A strong reversed move that uses $\alpha$ on the pair of values $\sum E_{k}$ and $\sum E_{k}^{2}$ is given by
$\sum E_{k-1}$ $=\sum E_{k}-\varepsilon\alpha$
$\sum E_{k-1}^{2}$ $=\sum E_{k}^{2}-2\sum E_{k}+(1-\varepsilon)^{2}\alpha^{2}.$
For integers $1\leq k\leq m$, let $\alpha_{k}$ denote the value of $\alpha$ used in the strong reversed move to get from $\left(\sum E_{k},\sum E_{k}^{2}\right)$ to $\left(\sum E_{k-1},\sum E_{k-1}^{2}\right)$. Then,
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\alpha_{1}+\varepsilon\alpha_{2}+\cdots+\varepsilon\alpha_{m}$
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}.$ (3)
The same process is a little more complicated for $\sum E_{m}^{2}$.
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
We can rewrite the $\sum E_{j}$ by plugging in (3).
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{j}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2\sum_{j=1}^{m}\left(\sum E_{0}+\varepsilon\sum_{i=1}^{j}\alpha_{i}\right)-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$\sum E_{m}^{2}$ $=\sum E_{0}^{2}+2m\sum E_{0}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $=\sum E_{0}^{2}+2m\sum E_{0}-\sum E_{m}^{2}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
We have $\sum E_{m}^{2}\geq\sum E_{m}$. Furthermore, due to (3), $\sum E_{m}\geq\varepsilon\sum_{j=1}^{m}\alpha_{j}$. We can use these two inequalities to replace $\sum E_{m}^{2}$.
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m-j+1)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}-\varepsilon\sum_{j=1}^{m}\alpha_{j}+2\varepsilon\sum_{j=1}^{m}(m)\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}\sum_{j=1}^{m}\alpha_{j}^{2}$ (4)
By the Cauchy-Schwarz Inequality, we have
$\left(\sum_{j=1}^{m}1\right)\left(\sum_{j=1}^{m}\alpha_{j}^{2}\right)$ $\geq\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
$\sum_{j=1}^{m}\alpha_{j}^{2}$ $\geq m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}.$
We can then plug this into (4):
$0$ $\leq\sum E_{0}^{2}+2m\sum E_{0}+\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}$
$0$ $\geq(1-\varepsilon)^{2}m^{-1}\left(\sum_{j=1}^{m}\alpha_{j}\right)^{2}-\varepsilon(2m-1)\sum_{j=1}^{m}\alpha_{j}-\left(\sum E_{0}^{2}+2m\sum E_{0}\right).$
We are now left with a quadratic in the sum $\sum_{j=1}^{m}\alpha_{j}$. Since the leading coefficient is positive and the quadratic is on the lesser side of the inequality, $\sum_{j=1}^{m}\alpha_{j}$ must be less than the larger root of this quadratic. This larger root is obtained through a tedious application of the quadratic formula:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\varepsilon(2m-1)+\sqrt{\varepsilon^{2}(2m-1)^{2}+4(1-\varepsilon)^{2}m^{-1}\left(\sum E_{0}^{2}+2m\sum E_{0}\right)}}{(1-\varepsilon)^{2}m^{-1}}.$ (5)
Lemma 16.
If $\varepsilon=\Omega\left(n^{-1/4}\right)$, then $\sum_{j=1}^{m}\alpha_{j}$ is bounded by $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
@@
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum E_{0}^{2}\right)+\Theta\left(\sum E_{0}\right)}}{\Theta\left(m^{-1}\right)}.$
Notice that we must have $\sum E_{0}^{2}\leq\left(\sum E_{0}\right)$. By Lemma 11, we can rewrite the previous inequality as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)+\Theta\left(m|S_{0}|\right)}}{\Theta\left(m^{-1}\right)}$
It is clear that the $\Theta\left(m|S_{0}|^{2}\right)$ term dominates the $\Theta\left(m|S_{0}|\right)$ term. We now have the following:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}.$ (6)
@@
$\Theta\left(\varepsilon^{2}m^{2}\right)$ dominates $\Theta\left(m|S_{0}|^{2}\right)$.
In this case, we can rewrite (6) as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon m)}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon m^{2}\right).$
Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$. This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
@@
In this case, we can rewrite (6) as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}m^{2})}+\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta\left(m|S_{0}|^{2}\right)}}{\Theta\left(m^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(m^{3/2}|S_{0}|\right).$
By Lemma 8, $\varepsilon=\Omega\left(n^{-1/4}\right)$ is equivalent to $\varepsilon=\Omega\left(m^{-1/4}\right)$.
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}m^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(m^{-1/4}\right)^{-2}m^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$
The result of Lemma 16 can be plugged into (3) with Lemma 11 to obtain a bound on $\sum E_{m}$. Since we are only interested in $\varepsilon=\Omega\left(n^{-1/4}\right)\geq\Theta\left(n^{-1/3}\right)$, we have
$\sum E_{m}$ $=\sum E_{0}+\varepsilon\sum_{j=1}^{m}\alpha_{j}$
$\sum E_{m}$ $=O\left(m|S_{0}|\right)+O\left(\varepsilon^{3}m^{2}|S_{0}|\right)$
$\sum E_{m}$ $=O\left(\varepsilon^{3}m^{2}|S_{0}|\right).$ (7)
By (7) and Lemma 10, the final number of bad edges in the graph must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. By construction, every vertex in $S_{0}$ begins with at least one bad edge, so we must begin with $\Omega\left(|S_{0}|\right)$ bad edges. We obtain the following upper bound on the price of uncertainty for consensus games when $\varepsilon=\Omega\left(n^{-1/4}\right)$:
$\frac{O\left(\varepsilon^{2}n^{2}|S_{0}|\right)}{\Omega\left(|S_{0}|\right)}=O\left(\varepsilon^{2}n^{2}\right).$

### v2

@@ 4.1 Proof of Improved Upper Bound
Suppose that $\varepsilon=\Omega(n^{-1/4})$. Then, $\sum_{i=1}^{m}z_{j}$ is $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$.
@@
$\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon m)+\sqrt{\Theta\left(\varepsilon^{2}m^{2}\right)+\Theta\left(m^{-1}\sum A_{0}^{2}\right)}}{\Theta\left(m^{-1}\right)}.$
By Lemma 7, we can replace every $m$ in the RHS with an $n$ to get
$\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+\Theta\left(n^{-1}\sum A_{0}^{2}\right)}}{\Theta\left(n^{-1}\right)}.$
By Eq. 2 and Proposition 11, $\sum A_{0}^{2}\leq(\sum A_{0})^{2}=O(n^{2}|S_{0}|^{2})$, so we can rewrite the previous inequality as follows:
$\sum_{j=1}^{m}z_{j}\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}.$ (10)
@@
$\Theta\left(\varepsilon^{2}n^{2}\right)$ dominates $\Theta\left(n|S_{0}|^{2}\right)$.
In this case, we can rewrite Eq. 10 as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\Theta(\varepsilon n)}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon n^{2}\right).$
Recall that $|S_{0}|=\Omega\left(\varepsilon^{-1}\right)$ because we must begin with that many bad edges. This means that $O\left(\varepsilon^{2}m^{2}|S_{0}|\right)$ must include $O\left(\varepsilon m^{2}\right)$, and we are done with this case.
@@
In this case, we can rewrite Eq. 10 as follows:
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{\Theta\left(\varepsilon^{2}n^{2}\right)+O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{\Theta(\varepsilon^{2}n^{2})}+\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $\leq\frac{\sqrt{O\left(n|S_{0}|^{2}\right)}}{\Theta\left(n^{-1}\right)}$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(n^{3/2}|S_{0}|\right).$
Since $\varepsilon=\Omega(n^{-1/4})$,
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\cdot\varepsilon^{-2}n^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}\left(n^{-1/4}\right)^{-2}n^{3/2}|S_{0}|\right)$
$\sum_{j=1}^{m}\alpha_{j}$ $=O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$
The result of Lemma 16 can be plugged into Eq. 7 with Proposition 11 to obtain a bound on $\sum E_{m}$. Since $\varepsilon=\Omega(n^{-1/4})$, we have
$\sum A_{m}$ $=\sum A_{0}+\varepsilon\sum_{i=1}^{m}z_{i}$
$\sum A_{m}$ $=O\left(n|S_{0}|\right)+O\left(\varepsilon^{3}n^{2}|S_{0}|\right)$
$\sum A_{m}$ $=O\left(\varepsilon^{3}n^{2}|S_{0}|\right).$ (11)
By Eq. 11 and Lemma 12, the final number of bad edges in the graph must be $O\left(\varepsilon^{2}n^{2}|S_{0}|\right)$. By construction, every vertex in $S_{0}$ begins with at least one bad edge, so we must begin with $\Omega\left(|S_{0}|\right)$ bad edges. We obtain the following upper bound on the price of uncertainty for consensus games when $\varepsilon=\Omega(n^{-1/4})$:
$\operatorname{PoU}(\varepsilon,\text{consensus})=\frac{O\left(\varepsilon^{2}n^{2}|S_{0}|\right)}{\Omega\left(|S_{0}|\right)}=O\left(\varepsilon^{2}n^{2}\right).$
@@
5 Conclusion
In this paper, we established tight asymptotic bounds on the worst-case increase in social cost experienced by a network under uncertainty. Our work leaves open some interesting future research directions. Firstly, we believe that the lower bound also holds in the more general case of arbitrary agent update ordering. This may more closely reflect the dynamics of real social networks, and we have a modified version of our construction in this paper that could potentially work for this new problem. Additionally, the study of adversarial influences to normal operations due to uncertainty or individual biases could also be extended to other game-theoretic network settings. While agents in our consensus game model have fixed relationships defined by graph edges, our model does not totally capture the fluidity of social networks due to the formation and destruction of individual connections. Incorporating uncertainty into more general models for social networks can provide greater insight into their behavior in the real world.